# Robust $M$-Estimation of Scatter Matrices via Precision Structure Shrinkage

### v1

@@ 5 Numerical experiments
In this section, we compare the finite-sample performance of the proposed estimator with those of the SCM, TME, LNSMI, and SSCM, as well as with $I_{p}$, using the Tyler-type weighting function w(s)=p/s(s>0)w(s)=p/s\;(s>0).
@@ 1. $\langle 1\rangle$
    Fix the data dimension $p$, specify the trace of the true scatter matrix $\Sigma\in\mathcal{S}_{++}^{p}$ as $\mathrm{Tr}[\Sigma]$, the largest contribution ratio $c_{1}\coloneqq\lambda_{1}(\Sigma)/\mathrm{Tr}[\Sigma]$, and the condition number $c_{2}\coloneqq\lambda_{1}(\Sigma)/\lambda_{p}(\Sigma)$, e.g., $\mathrm{Tr}[\Sigma]=p$, $c_{1}=0.30$, $c_{2}=50$, and generate $\Sigma=\Sigma(c_{1},c_{2})$ that satisfies these specifications. Since all methods compared here possess orthogonal equivariance as in Subsection 3.4, we may choose the eigenvectors of $\Sigma$ as the standard basis.
@@ 2. $\langle 2\rangle$
    Generation of main-body data
    Specify all sample sizes $N$ and outlier ratios $\xi$, e.g., $N=100$, $\xi=0.03$, and set the number of outliers as $m=\lceil N\xi\rceil$. Generate the main body of data $\bm{z}_{i}$ for $i\in[n]$, where $n\coloneqq N-m$ from the multivariate normal distribution with mean vector 0 and scatter matrix $\Sigma$.
@@ 3. $\langle 3\rangle$
    Generate the outliers $\bm{z}_{i}$ for $i\in[N]\setminus[n]$ so that their squared Mahalanobis distances are a constant multiple $k$, e.g., $k=10$, of the sample mean for the Mahalanobis distances of the main-body data points
@@
      Obtain $m$ points from random directions, shown in Figure 2-(a);
@@
      Obtain $m$ points from a single random direction, as a cluster, shown in Figure 2-(b). Specifically, outliers are randomly generated from $\mathcal{N}(\bm{\mu}_{\mathrm{o}},0.01I_{p})$, where $\bm{\mu}_{\mathrm{o}}$ is given by the method $\langle 3\rangle$.
Figure 1: An illustration of the distribution of outliers for the case $p=2$, with $c_{1}=0.6$, $c_{2}=1000$, $N=10000$, $\xi=0.1$, and $k=10$. Data points belonging to the main body are plotted as black circles, whereas outliers are represented by red crosses.
Figure 2: An illustration of the outlier generation methods for the case $p=2$, with $c_{1}=0.6$, $c_{2}=1000$, $N=100$, $\xi=0.03$, and $k=10$. The left panel illustrates unclustered outliers, while the right panel illustrates clustered outliers. Data points belonging to the main body are plotted as black circles, whereas outliers are represented by red crosses.
By varying the combinations of (i) data generation parameters $(p,\mathrm{Tr}[\Sigma],c_{1},c_{2},N,\xi,k)$, (ii) the generating distribution of the main body data, and (iii) the outlier configurations (unclustered outliers or clustered outliers), the performance of SCM, TME, LNSMI, SSCM, $I_{p}$, and the proposed Tyler-type method is evaluated under each condition. The performance of each estimator is assessed using the RMSE computed over 1000 trials for each setting. In addition, the shrinkage coefficient of LNSMI is determined based on Theorem 2 in [5].
First, Subsection 5.3 examines how the estimation error changes with varying dimensions of the random vector. These results demonstrate that TME and LNSMI fail to perform properly under certain conditions. Next, Subsection 5.4 examines how the estimation error changes when the shrinkage coefficient $\alpha$ is varied. This analysis reveals the behavior of LNSMI and the proposed Tyler-type estimator as estimators, regardless of whether the shrinkage coefficient is successfully computed. In this analysis, we examine the effect of the shrinkage coefficient on the estimator.
@@
Once the largest contribution ratio $c_{1}$, the condition number $c_{2}$, and the trace $\mathrm{Tr}[\Sigma]$ are specified, the values of $\lambda_{1}(\Sigma)$ and $\lambda_{p}(\Sigma)$ are uniquely determined. Details of how $\lambda_{2}(\Sigma),\ldots,\lambda_{p-1}(\Sigma)$ are generated are provided in Appendix C.
@@
The initial values for TME, LNSMI, and the proposed method are all set to the trace-normalized SCM. For TME and LNSMI, the stopping criterion is defined as
$\left\|V_{k+1}-V_{k}\right\|_{\textnormal{F}}<\epsilon,$
with $\epsilon=10^{-6}$. For the proposed method, among several stopping criteria examined, the following condition yielded the best performance and is therefore adopted:
$\min\left\{\left\|\Omega_{k+1}-\Omega_{k}\right\|_{\textnormal{F}},\left\|V_{k+1}-V_{k}\right\|_{\textnormal{F}}\right\}<\epsilon\text{.}$
@@ 5.2 Penalty coefficient computation for the proposed method
Following [4, 5, 23], we determine the shrinkage coefficient $\hat{\alpha}$ by considering the optimization problem
$\underset{\alpha\in\mathbb{R}}{\operatorname{arg\,min}}\ \textnormal{E}\left[\left\|(1-\alpha)C_{\textnormal{P}}^{-1}+\alpha I_{p}-\Omega\right\|_{\textnormal{F}}^{2}\right]\text{, where}\;C_{\textnormal{P}}\coloneqq\frac{p}{n}\sum_{i=1}^{N}\frac{\bm{x}_{i}\bm{x}_{i}^{\top}}{\bm{x}_{i}^{\top}\Omega\bm{x}_{i}}\text{.}$
@@
$\alpha_{\textnormal{o}}\coloneqq\frac{\mathrm{Tr}[\Omega]-\mathrm{Tr}\left[\textnormal{E}\left[C_{\textnormal{P}}^{-1}\right]\right]-\mathrm{Tr}\left[\textnormal{E}\left[C_{\textnormal{P}}^{-1}\right]\Omega\right]+\textnormal{E}\left[\left\|C_{\textnormal{P}}^{-1}\right\|_{\textnormal{F}}^{2}\right]}{p-2\mathrm{Tr}\left[\textnormal{E}\left[C_{\textnormal{P}}^{-1}\right]\right]+\textnormal{E}\left[\left\|C_{\textnormal{P}}^{-1}\right\|_{\textnormal{F}}^{2}\right]}\text{.}$ (14)
@@
$\frac{\mathrm{Tr}[M_{2}]-\mathrm{Tr}\left[\textnormal{E}\left[\hat{C}_{\textnormal{P}}^{-1}\right]\right]-\mathrm{Tr}\left[\textnormal{E}\left[\hat{C}_{\textnormal{P}}^{-1}\right]M_{2}\right]+\textnormal{E}\left[\left\|\hat{C}_{\textnormal{P}}^{-1}\right\|_{\textnormal{F}}^{2}\right]}{p-2\mathrm{Tr}\left[\textnormal{E}\left[\hat{C}_{\textnormal{P}}^{-1}\right]\right]+\textnormal{E}\left[\left\|\hat{C}_{\textnormal{P}}^{-1}\right\|_{\textnormal{F}}^{2}\right]}$
as $\hat{\alpha}$. However, because it is often difficult to express $\textnormal{E}[\hat{C}_{\textnormal{P}}^{-1}]$ and $\textnormal{E}[\|\hat{C}_{\textnormal{P}}^{-1}\|_{\textnormal{F}}^{2}]$ in closed form, we further propose to replace these expectations by their numerical approximations obtained via the bootstrap, and to use the resulting plug-in value as $\hat{\alpha}$. Specifically, by substituting the bootstrap approximations $m_{1}$ and $m_{2}$ for $\textnormal{E}[\hat{C}_{\textnormal{P}}^{-1}]$ and $\textnormal{E}[\|\hat{C}_{\textnormal{P}}^{-1}\|_{\textnormal{F}}^{2}]$, respectively, we propose the shrinkage coefficient
@@
When approximating $\textnormal{E}[\hat{C}_{\textnormal{P}}^{-1}]$ and $\textnormal{E}[\|\hat{C}_{\textnormal{P}}^{-1}\|_{\textnormal{F}}^{2}]$ via the bootstrap method, it is necessary to resample as many distinct data points as possible, up to a maximum of $N$, in order to avoid rank deficiency in $\hat{C}_{\textnormal{P}}$. Specifically, when the sample size is $N=100$, we recommend performing approximately 370 resampling iterations. The justification for this choice is provided in Appendix B.
@@ 5.3 Performance across different dimensions
Figure 3: Plots of the RMSE values of each estimator for different data dimensions $p$, with lines connecting the plotted points. The horizontal axis represents the dimension $p$ $(=2,\ldots,55)$, and the vertical axis represents the RMSE. The other settings are $N=100$, $\xi=0.03$, and $k=10$. Figure (a, c, e) show the results for unclustered outliers, while Figure (b, d, f) correspond to clustered outliers. In Figure (a, c, e), the curves representing TME, LNSMI, SSCM, and the proposed method nearly overlap. Furthermore, in Figure (d, f), the curves representing SSCM and the proposed method nearly overlap.
@@
Figure 3-(a, c, e) shows the results for the case of unclustered outliers. Specifically, Figure 3-(a) corresponds to $(c_{1},c_{2})=(0.6,1000)$, Figure 3-(c) to $(c_{1},c_{2})=(0.3,50)$, and Figure 3-(e) to $(c_{1},c_{2})=(1/p,1)$, that is, the case where $\Sigma=I_{p}$.
The curves representing the performance of TME, LNSMI, SSCM, and the proposed Tyler-type method nearly overlap. However, in Figure 3-(a), where the true covariance structure deviates from the identity structure $I_{p}$, SSCM exhibits relatively poor estimation accuracy. On the other hand, the behavior of all estimators remains highly stable, and in particular, TME, LNSMI, SSCM, and the proposed method show better estimation accuracy than SCM.
@@
Figure 3-(b, d, f) shows the results for the case of clustered outliers. Specifically, Figure 3-(b) corresponds to $(c_{1},c_{2})=(0.6,1000)$, Figure 3-(d) to $(c_{1},c_{2})=(0.3,50)$, and Figure 3-(f) to $(c_{1},c_{2})=(1/p,1)$.
It is observed that when the outlier ratio is $\xi=0.03$, the performance of TME and LNSMI deteriorates markedly once the dimension exceeds $p=35$, becoming even worse than that of SCM. In contrast, SSCM and the proposed method successfully suppress the severe degradation seen in TME and LNSMI. These phenomena are consistently observed under various settings with different values of $\mathrm{Tr}[\Sigma]$, $N$, $\xi$, and $k$. It is known that the asymptotic breakdown point of TME lies between $1/(p+1)$ and $1/p$, and the observed behavior of TME conforms to this property [9].
@@ 5.4 Sensitivity to shrinkage coefficients
We set the shrinkage coefficient $\alpha\in[0,1]$ to various values for LNSMI and the proposed Tyler-type method, and compare the behavior of each estimator. We set $p\in{25,50}$, $\mathrm{Tr}[\Sigma]=p$, $N=100$, $\xi=0.03$, and $k=10$. Figure 4 presents the results for $p=25$, and Figure 5 shows those for $p=50$. In Figures 4 and 5, panels (a, c, e) correspond to the case of unclustered outliers, while panels (b, d, f) correspond to the case of clustered outliers. Moreover, in Figures 4 and 5, panels (a, b) show the results for $(c_{1},c_{2})=(0.6,1000)$, panels (c, d) for $(c_{1},c_{2})=(0.3,50)$, and panels (e, f) for $(c_{1},c_{2})=(1/p,1)$.
Since SCM, TME, and SSCM do not involve shrinkage, their corresponding results remain constant. For LNSMI, when $\alpha=1$, the output becomes $I_{p}$. Therefore, in Figure 4-(e), Figure 4-(f), Figure 5-(e), and Figure 5-(f), where the true covariance structure is $I_{p}$, the vertical axis value of LNSMI becomes 0 at $\alpha=1$. In contrast, for the proposed method, the result coincides with SSCM when $\alpha=1$. The estimation errors of TME and those of LNSMI and the proposed Tyler-type method show very similar behavior when $\alpha\approx 0$. In fact, when $\alpha\approx 0$, both LNSMI and the proposed method operate in a manner similar to TME. In particular, LNSMI and the proposed method theoretically coincide when $\alpha=0$.
We now provide an overview of the behavior of the estimators as the shrinkage coefficient $\alpha$ is varied. Note that the meaning of nonzero values of $\alpha$ differs between LNSMI and the proposed method, so the results of the two methods cannot be directly compared when $\alpha\neq 0$.
First, comparing the points at which each estimator attains its smallest estimation error over the range $\alpha\in[0,1]$, we see that the proposed method achieves accuracy comparable to the other estimators in many settings. Conversely, in the presence of clustered outliers, LNSMI can also attain estimation accuracy similar to that of the proposed method if an appropriate value of $\alpha$ is chosen. However, except for the cases in Figure 4-(e), Figure 4-(f), Figure 5-(e), and Figure 5-(f), where the true covariance structure is $I_{p}$, the curve representing the proposed method generally lies below that of LNSMI. This observation suggests that the proposed method is robust with respect to the choice of $\alpha$.

### v2

@@ 5 Numerical experiments
In this section, we compare the finite sample performance of the proposed estimator, which uses the Tyler type weighting function w(s)=p/s(s>0)w(s)=p/s\,(s>0), with that of the SCM, TME, LNSMI, SSCM, and $I_{p}$.
@@ 1. $\langle 1\rangle$
    Fix the data dimension $p$, specify the trace $\mathrm{Tr}[\Sigma]$ of the true scatter matrix $\Sigma\in\mathcal{S}_{++}^{p}$, the largest contribution ratio $c_{1}\coloneqq\lambda_{1}(\Sigma)/\mathrm{Tr}[\Sigma]$, and the condition number $c_{2}\coloneqq\lambda_{1}(\Sigma)/\lambda_{p}(\Sigma)$, e.g., $\mathrm{Tr}[\Sigma]=p$, $c_{1}=0.30$, $c_{2}=50$, and generate $\Sigma=\Sigma(c_{1},c_{2})$ that satisfies these specifications. Since all methods compared here possess orthogonal equivariance as discussed in Subsection 3.4, we may choose the eigenvectors of $\Sigma$ as the standard basis.
@@ 2. $\langle 2\rangle$
    Generation of main body data
    Specify all sample sizes $N$ and outlier ratios $\xi$, e.g., $N=100$, $\xi=0.03$, and set the number of outliers as $m=\lceil N\xi\rceil$. Generate the main body data $\bm{z}_{i}$ for $i\in[n]$, where $n\coloneqq N-m$ from the multivariate normal distribution with mean vector $\bm{0}$ and scatter matrix $\Sigma$.
@@ 3. $\langle 3\rangle$
    Generate the outliers $\bm{z}_{i}$ for $i\in[N]\setminus[n]$ so that their squared Mahalanobis distances equal a constant multiple $k$, e.g., $k=10$, of the sample mean of the squared Mahalanobis distances of the main body data points
@@
      We obtain $m$ points from random directions, shown in Figure 2-(a);
@@
      We obtain $m$ points from a single random direction, as a cluster, shown in Figure 2-(b). Specifically, outliers are randomly generated from $\mathcal{N}_{p}(\bm{\mu}_{\mathrm{o}},0.01I_{p})$, where $\bm{\mu}_{\mathrm{o}}$ is given by the method $\langle 3\rangle$.
Figure 1: An illustration of the distribution of outliers for the case $p=2$, with $c_{1}=0.6$, $c_{2}=1000$, $N=10000$, $\xi=0.1$, and $k=10$. Data points belonging to the main body are plotted as black circles, and outliers are represented by red crosses.
Figure 2: An illustration of the outlier generation methods for the case $p=2$, with $c_{1}=0.6$, $c_{2}=1000$, $N=100$, $\xi=0.03$, and $k=10$. The left panel illustrates unclustered outliers, while the right panel illustrates clustered outliers. Data points belonging to the main body are plotted as black circles, and outliers are represented by red crosses.
The performance of SCM, TME, LNSMI, SSCM, $I_{p}$, and the proposed Tyler type estimator is evaluated under each condition by varying
  • (i)
    data generation parameters $(p,\mathrm{Tr}[\Sigma],c_{1},c_{2},N,\xi,k)$,
  • (ii)
    the generating distribution of the main body data,
  • (iii)
    the outlier configurations (unclustered outliers or clustered outliers).
The performance of each estimator is assessed using the RMSE computed over 1000 trials for each setting. In addition, the shrinkage coefficient of LNSMI is determined based on Theorem 2 in [5].
First, Subsection 5.3 examines how the estimation error changes with varying dimensions of the random vector. These results demonstrate that TME and LNSMI fail to perform properly under certain conditions. Next, Subsection 5.4 examines how the estimation error changes when the shrinkage coefficient $\alpha$ is varied. This analysis reveals the behavior of LNSMI and the proposed Tyler type estimator, regardless of whether the shrinkage coefficient is successfully computed. In this analysis, we examine the effect of the shrinkage coefficient on the estimator.
@@
Once the largest contribution ratio $c_{1}$, the condition number $c_{2}$, and the trace $\mathrm{Tr}[\Sigma]$ are specified, the largest and smallest eigenvalues $\lambda_{1}(\Sigma)$ and $\lambda_{p}(\Sigma)$ are uniquely determined. Details on the generation of $\lambda_{2}(\Sigma),\ldots,\lambda_{p-1}(\Sigma)$ can be found in Appendix C.
@@
The initial values for TME, LNSMI, and the proposed method are set to the trace normalized SCM. For TME and LNSMI, the stopping criterion is defined as
$\left\|V_{k+1}-V_{k}\right\|_{\mathrm{F}}<\epsilon,$
with $\epsilon=10^{-6}$. For the proposed method, among several stopping criteria examined, the following stopping criterion yielded the best performance and is therefore adopted:
$\min\left\{\left\|\Omega_{k+1}-\Omega_{k}\right\|_{\mathrm{F}},\left\|V_{k+1}-V_{k}\right\|_{\mathrm{F}}\right\}<\epsilon\text{.}$
@@ 5.2 Penalty coefficient computation for the proposed method
Following [4, 5, 18, 22], we determine the shrinkage coefficient $\hat{\alpha}$ by considering the optimization problem
$\underset{\alpha\in\mathbb{R}}{\operatorname{arg\,min}}\ \mathrm{E}\left[\left\|(1-\alpha)C_{\textnormal{P}}^{-1}+\alpha I_{p}-\Omega\right\|_{\mathrm{F}}^{2}\right]\text{, where}\;C_{\textnormal{P}}\coloneqq\frac{p}{N}\sum_{i=1}^{N}\frac{\bm{x}_{i}\bm{x}_{i}^{\top}}{\bm{x}_{i}^{\top}\Omega\bm{x}_{i}}\text{.}$
@@
$\alpha_{\textnormal{o}}\coloneqq\frac{\mathrm{Tr}[\Omega]-\mathrm{Tr}\left[\mathrm{E}\left[C_{\textnormal{P}}^{-1}\right]\right]-\mathrm{Tr}\left[\mathrm{E}\left[C_{\textnormal{P}}^{-1}\right]\Omega\right]+\mathrm{E}\left[\left\|C_{\textnormal{P}}^{-1}\right\|_{\mathrm{F}}^{2}\right]}{p-2\mathrm{Tr}\left[\mathrm{E}\left[C_{\textnormal{P}}^{-1}\right]\right]+\mathrm{E}\left[\left\|C_{\textnormal{P}}^{-1}\right\|_{\mathrm{F}}^{2}\right]}\text{.}$ (14)
@@
$\frac{\mathrm{Tr}[M_{2}]-\mathrm{Tr}\left[\mathrm{E}\left[\hat{C}_{\textnormal{P}}^{-1}\right]\right]-\mathrm{Tr}\left[\mathrm{E}\left[\hat{C}_{\textnormal{P}}^{-1}\right]M_{2}\right]+\mathrm{E}\left[\left\|\hat{C}_{\textnormal{P}}^{-1}\right\|_{\mathrm{F}}^{2}\right]}{p-2\mathrm{Tr}\left[\mathrm{E}\left[\hat{C}_{\textnormal{P}}^{-1}\right]\right]+\mathrm{E}\left[\left\|\hat{C}_{\textnormal{P}}^{-1}\right\|_{\mathrm{F}}^{2}\right]}$
as $\hat{\alpha}$. However, because it is often difficult to express $\mathrm{E}[\hat{C}_{\textnormal{P}}^{-1}]$ and $\mathrm{E}[\|\hat{C}_{\textnormal{P}}^{-1}\|_{\mathrm{F}}^{2}]$ in closed form, we therefore replace these expectations by their numerical approximations obtained via the bootstrap, and use the resulting plug-in value as $\hat{\alpha}$. Specifically, by substituting the bootstrap approximations $m_{1}$ and $m_{2}$ for $\mathrm{E}[\hat{C}_{\textnormal{P}}^{-1}]$ and $\mathrm{E}[\|\hat{C}_{\textnormal{P}}^{-1}\|_{\mathrm{F}}^{2}]$, respectively, we propose the shrinkage coefficient
@@
When approximating $\mathrm{E}[\hat{C}_{\textnormal{P}}^{-1}]$ and $\mathrm{E}[\|\hat{C}_{\textnormal{P}}^{-1}\|_{\mathrm{F}}^{2}]$ via the bootstrap method, it is necessary to resample as many distinct data points as possible, up to a maximum of $N$, in order to avoid rank deficiency in $\hat{C}_{\textnormal{P}}$. Specifically, when the sample size is $N=100$, we recommend performing approximately 370 resampling iterations. The justification for this choice is provided in Appendix B.
@@ 5.3 Performance across different dimensions
Figure 3: Plots of the RMSE values of each estimator for different data dimensions $p$, with lines connecting the plotted points. The horizontal axis represents the dimension $p$ $(=2,\ldots,55)$, and the vertical axis represents the RMSE. The other settings are $N=100$, $\xi=0.03$, and $k=10$. Figure (a, c, e) show the results for unclustered outliers, while Figure (b, d, f) correspond to clustered outliers. In Figure (a, c, e), the curves representing TME, LNSMI, SSCM, and the proposed Tyler type estimator nearly overlap. Furthermore, in Figure (d, f), the curves representing SSCM and the proposed Tyler type estimator nearly overlap.
@@
Figure 3-(a, c, e) shows the results for the case of unclustered outliers. Specifically, Figure 3-(a), (c), and (e) correspond to $(c_{1},c_{2})=(0.6,1000)$, $(0.3,50)$, and $(1/p,1)$, respectively, where the last case corresponds to $\Sigma=I_{p}$.
The curves representing the performance of TME, LNSMI, SSCM, and the proposed Tyler type estimator nearly overlap with one another. However, in Figure 3-(a), where the true covariance structure deviates from $I_{p}$, SSCM exhibits relatively poor estimation accuracy. On the other hand, the behavior of all estimators remains highly stable, and in particular, TME, LNSMI, SSCM, and the proposed Tyler type estimator show better estimation accuracy than SCM.
@@
Figure 3-(b, d, f) shows the results for the case of clustered outliers. Specifically, Figure 3-(b), (d), and (f) correspond to $(c_{1},c_{2})=(0.6,1000)$, $(0.3,50)$, and $(1/p,1)$, respectively.
We observe that when the outlier ratio is $\xi=0.03$, the performance of TME and LNSMI deteriorates markedly once the dimension exceeds $p=35$, becoming even worse than that of SCM. In contrast, SSCM and the proposed Tyler type estimator successfully suppress the severe degradation seen in TME and LNSMI. These phenomena are consistently observed under various settings with different values of $\mathrm{Tr}[\Sigma]$, $N$, $\xi$, and $k$. It is known that the asymptotic breakdown point of TME lies between $1/(p+1)$ and $1/p$, and the observed behavior of TME conforms to this property [8].
@@ 5.4 Sensitivity to shrinkage coefficients
To examine the effect of the shrinkage coefficient, we set $\alpha\in[0,1]$ to various values for LNSMI and the proposed Tyler type estimator and compare the behavior of each estimator. We set $p\in\{25,50\}$, $\mathrm{Tr}[\Sigma]=p$, $N=100$, $\xi=0.03$, and $k=10$. Figure 4 presents the results for $p=25$, and Figure 5 shows those for $p=50$. In Figures 4 and 5, panels (a, c, e) correspond to the case of unclustered outliers, while panels (b, d, f) correspond to the case of clustered outliers. Moreover, in Figures 4 and 5, panels (a, b), (c, d), and (e, f) correspond to $(c_{1},c_{2})=(0.6,1000)$, $(0.3,50)$, and $(1/p,1)$, respectively.
Since SCM, TME, and SSCM do not involve shrinkage, their corresponding results remain constant. For LNSMI, when $\alpha=1$, the output becomes $I_{p}$. Therefore, in Figure 4-(e, f) and Figure 5-(e, f), where the true covariance structure is $I_{p}$, the RMSE value of LNSMI becomes 0 at $\alpha=1$. In contrast, for the proposed Tyler type estimator, the result coincides with SSCM when $\alpha=1$. The estimation errors of TME and those of LNSMI and the proposed Tyler type estimator show very similar behavior when $\alpha\approx 0$. Indeed, when $\alpha\approx 0$, both LNSMI and the proposed Tyler type estimator operate in a manner similar to TME. In particular, LNSMI and the proposed Tyler type estimator theoretically coincide when $\alpha=0$.
Below, we provide an overview of the behavior of the estimators as the shrinkage coefficient $\alpha$ is varied. Note that the meaning of nonzero values of $\alpha$ differs between LNSMI and the proposed method, so the results of the two methods cannot be directly compared when $\alpha\neq 0$.
First, comparing the points at which each estimator attains its smallest estimation error over the range $\alpha\in[0,1]$, we see that the proposed Tyler type estimator achieves accuracy comparable to the other estimators in many settings. On the other hand, in the presence of clustered outliers, LNSMI can also attain estimation accuracy similar to that of the proposed Tyler type estimator if an appropriate value of $\alpha$ is chosen. However, except for the cases in Figure 4-(e, f) and Figure 5-(e, f), where the true covariance structure is $I_{p}$, the curve representing the proposed Tyler type estimator generally lies below that of LNSMI. This observation suggests that the proposed Tyler type estimator is robust with respect to the choice of $\alpha$.